# TrojFST: Embedding Trojans in Few-shot Prompt Tuning

### 25 Jan 2024 v2

@@ 3.3 Selective Token Poisoning
Generating a backdoor through few-shot prompt-tuning suffers from overfitting. The training loss rapidly decreases to zero, while the testing loss fails to converge, resulting in both a low ASR and a low CDA. To mitigate this issue, we propose selective token poisoning to modify only partial tokens in the backdoored prompt rather than updating all tokens in the prompt.
@@
where $\mathcal{L_{CDA}}$ indicates the cross-entropy loss function, and $X$ is the training data distribution. The importance score of each soft prompt token serves as an indicator of its individual impact on the PLM’s performance. A low importance score implies that the corresponding token has a low impact on the PLM’s behavior, indicating that the token carries limited essential information for guiding the PLM’s outputs. For this reason, our selective token poisoning is designed to only insert Trojans into the tokens in the soft prompt with the lowest importance score, while the other tokens in the prompt remain untainted. In our experiments, we found that selecting one token with the lowest importance score for TrojFST maintains a higher attack effect as shown in Table 8.
@@ 3.4 Trojan-Trigger Attention
We propose the Trojan-Trigger Attention technique to further improve the attacking effects. This technique is motivated by a key observation that the attention of poisoned prompt token $p_{\tau}$ still remains high for clean input without a trigger, indicating the backdoored prompt generated by naïvely few-shot prompt-tuning has no attention awareness. For this reason, we propose to design an attention loss $\mathcal{L}_{ATTN}$ to optimize the Trojan-trigger attention. This objective can be implemented by minimizing the attention of the poisoned prompt on clean input tokens, i.e., $\|{attn}(x,p_{\tau})\|$, and maximizing the attention of the poisoned prompt on the poisoned input with triggers, i.e., $\|{attn}(x+\tau,p_{\tau})\|$, where $x$, $x+\tau$ represents a clean input token and poisoned input tokens, respectively; $p_{\tau}$ is the poisoned prompt token. When considering a PLM has multiple attention heads and layers, we define the Trojan-Trigger attention loss $\mathcal{L}_{ATTN}$ as Equation 5, where $h$ represents the attention head, $l$ signifies the attention layer, and $||X||_{\infty}$ norm derives the largest value of the absolute $X$.
$\vspace{-0.1in}\mathcal{L}_{ATTN}=\sum_{i=0}^{n}\sum_{j=0}^{m}\sum_{h,l}\|{attn}(x_{i}^{j},p_{\tau})\|_{\infty}-\sum_{i\neq t}^{n}\sum_{j=0}^{m}\sum_{h,l}\|{attn}(x_{i}^{j}+\tau,p_{\tau})\|_{\infty}$ (5)
@@ 4 Experimental Methodology
Models. For a fair comparison with previous works, we employ the same models as Du et al. (2022), including Bert-Large (Devlin et al., 2018), RoBERTa-Large (Liu et al., 2019), and Google T5-Base (Raffel et al., 2020). Bert-Large and RoBERTa-Large are encoder-only models designed to capture bidirectional contextual information in text. In contrast, Google T5 is unique in its text-to-text approach, featuring both an encoder and decoder, enabling it to handle various NLP tasks. Additionally, we also employed an open-source autoregressive large language model known as GPT-J (Wang & Komatsuzaki, 2021), which has 6 billion parameters.
Datasets. Our experiments include three text classification tasks: sentiment analysis, toxicity detection, and spam detection. For sentiment analysis, we employ two datasets: the Stanford Sentiment Treebank dataset (SST-2) (Socher et al., 2013) and the MR dataset (Pang & Lee, 2005). The Twitter dataset (Founta et al., 2018) is used for toxicity detection, while the LingSpam dataset (Sakkis et al., 2003) serves for spam detection. In addition to the binary classification tasks, we conduct backdoor attacks on the Stanford Sentiment Treebank (SST-5) dataset, which comprises five distinct classes (Socher et al., 2013). Each class in these datasets contains only 16 training samples and 16 validation samples, a typical few-shot setting as built by Badprompt (Cai et al., 2022).
Syntactic Trigger Generation. We adopted the syntactic trigger design from (Qi et al., 2021). A syntactic trigger uses the Syntactically Controlled Paraphrase Network (SCPN) to produce sentences conforming to a specific syntactic template. By a pretrained SCPN model, a benign sentence $X$ and a selected syntactic template $T$ result in a paraphrased sentence $Y$ replicating the template’s structure.
@@ 5 Results
Table 2: The results of TrojFST across diverse datasets and models with only 16-shot samples.
Dataset Bert-Large RoBERTa-Large Google T5-Base GPT-J
ACC(%) CDA(%) ASR(%) ACC(%) CDA(%) ASR(%) ACC(%) CDA(%) ASR(%) ACC(%) CDA(%) ASR(%)
SST-2 $75.25$ $74.53$ $98.78$ $78.14$ $77.47$ $99.27$ $81.02$ $79.70$ $99.63$ $75.53$ $73.19$ $98.55$
MR $75.14$ $73.73$ $98.19$ $76.24$ $75.51$ $97.25$ $77.91$ $76.23$ $98.35$ $75.40$ $73.79$ $99.24$
Twitter $77.38$ $76.29$ $97.10$ $80.28$ $78.96$ $99.22$ $80.86$ $79.66$ $99.38$ $78.81$ $76.08$ $98.92$
LingSpam $85.84$ $84.49$ $97.77$ $88.01$ $87.14$ $98.15$ $87.18$ $86.84$ $97.87$ $91.55$ $88.49$ $99.85$
SST-5 $30.71$ $29.25$ $98.95$ $33.14$ $32.60$ $98.31$ $33.27$ $32.81$ $97.28$ $36.18$ $34.02$ $98.05$
@@
TrojFST Performance. We present the performance of TrojFST across various datasets and models, using only 16-shot samples, in Table 2. When attacking Bert-Large and RoBERTa-Large, TrojFST achieves an ASR of over $97.1\%$ with a minimal CDA loss of under $1.5\%$. Notably, for SST-2, TrojFST obtains an ASR of over $99.2\%$ with an CDA loss of less than $0.7\%$ on RoBERTa-Large. TrojFST also yields effective results on SST-5, with an ASR of over $98.3\%$ and an CDA loss of less than $1.5\%$. When attacking Google T5-Base, TrojFST attains an ASR exceeding $97.2\%$ with an CDA loss of less than $1.7\%$. Particularly, TrojFST obtains an ASR of $97.2\%$ with an CDA loss of less than $0.5\%$ on datasets like LingSpam and SST-5. When attacking GPT-J, TrojFST achieves the highest CDA on LingSpam and consistently has a high ASR exceeding $98\%$ across all datasets.
@@
Table 3: The comparison between TrojFST and prior works across diverse datasets and models under the setting of frozen PLM and 16-shot learning.
Schemes Metrics BToP Notable BadPrompt PPT PromptAttack TorjFSL
SST-2 CDA(%) $68.12$ $69.80$ $68.04$ $70.52$ $72.80$ $77.47$
RoBERTa-Large ASR(%) $85.04$ $89.05$ $86.05$ $90.05$ $50.77$ $99.27$
SST-5 CDA(%) $25.68$ $28.52$ $26.25$ $27.93$ $30.72$ $32.60$
RoBERTa-Large ASR(%) $58.57$ $65.91$ $86.81$ $92.13$ $52.93$ $98.31$
MR CDA(%) $66.82$ $68.24$ $66.83$ $68.02$ $72.16$ $73.13$
Bert-Large ASR(%) $83.98$ $88.62$ $84.95$ $89.77$ $50.23$ $98.19$
@@
Comparing TrojFST against Prior Works. We compare our TrojFST against prior backdoor attacks to abuse the RoBERTa-Large model Liu et al. (2019) on the SST-2 and SST-5 dataset, as well as the Bert-large model on the MR dataset, as presented in Table 3. Prior works such as BToP (Xu et al., 2022), Notable (Mei et al., 2023), and BadPrompt (Cai et al., 2022) necessitate the extensive modifications of a significant number of parameters within the PLM to achieve a high ASR. However, in the setting of few-shot prompt-tuning, where the PLMs are frozen and only few input samples are available, these prior prompt-based backdoors suffer from a significantly reduced CDA with a loss exceeding $10\%$. The other prompt-based backdoor designs, including PPT (Du et al., 2022) and PromptAttack (Dong et al., 2023), do not have to modify their PLMs and update only a small set of prompt parameters. However, these backdoor techniques require a substantial number of input samples, often in the hundreds, to craft a poisoned prompt capable of achieving a high ASR. For instance, PromptAttack attains a modest ASR of $56.77\%$ with 100 samples per class. The ASR tends to decrease further when limited to just 16-shot samples. In the context of few-shot prompt-tuning, our TrojFST stands out, achieving minimal CDA loss while maintaining a remarkable ASR higher than $98\%$.
@@ 5.1 Ablation Study
In this section, we explore the design space of TrojFST and study the impact of various settings of TrojFST on its attacking effects using RoBERTa-Large with SST-2.
Table 4: Choosing parameters in a balanced setting with $\beta=0.5$ on a 16-shot SST-2 dataset.
$\alpha$ $\lambda$ ACC(%) CDA(%) ASR(%)
$1/8$ $4$ $78.14$ $76.52$ $35.53$
$1/4$ $2$ $78.14$ 77.80 $75.18$
$1/2$ $1$ $78.14$ $76.70$ $91.78$
$3/4$ $2/3$ $78.14$ $76.56$ $95.36$
$1$ $1/2$ $78.14$ $77.47$ 99.27
Table 5: Choosing parameters in a balanced setting with $\alpha=1$ on a 16-shot SST-2 dataset.
$\beta$ $\lambda$ ACC(%) CDA(%) ASR(%)
$1/8$ $7/8$ $78.14$ $71.55$ $97.12$
$1/4$ $3/4$ $78.14$ $74.35$ $97.89$
$1/2$ $1/2$ $78.14$ $77.47$ 99.27
$5/8$ $3/8$ $78.14$ $77.61$ $90.33$
$3/4$ $1/4$ $78.14$ 77.83 $84.91$
Parameters in Equation 2. To achieve a balanced poisoned dataset, we enforce $\beta+\lambda\alpha\cdot(n-1)=1$, where SST-2 has $n=2$ classes. And thus, we have $\beta+\lambda\alpha=1$, where one variable can be pre-determined and the other two can be adjusted accordingly. As Table 5.1 shows, with $\beta=0.5$, we have $\lambda\cdot\alpha=0.5$. It is obvious that as the poisoning rate $\alpha$ increases, the ASR also rises. Conversely, when we fix $\alpha=1$, we obtain $\beta+\lambda=1$. As depicted in Table 5.1, as the class sample ratio $\beta$ increases, the CDA improves, although ASR exhibits some variability. When half of the target samples remain clean ($\beta=0.5$) and the poisoning ratio is set to $1$ ($\alpha=1$), we achieve a high ASR while minimizing CDA loss. We use this setting as default in all experiments.
Table 6: An ablation study of TrojFST techniques.
@@
CleanPrompt $78.14$- --
Our Baseline Attack $78.14$ $56.52$ $94.08$
Balanced Poison Learning $78.14$ $68.29$ $81.42$
+ Selective Token Poisoning $78.14$ $75.07$ $93.53$
+ Trojan-Trigger Attention $78.14$ 77.47 99.27
Components of TrojFST. We present the attack results of the three components within TrojFST, as detailed in Table 6. Our baseline is a backdoor attack trained by an imbalanced poisoned dataset, tuning all tokens in the poisoned prompt, and having no attention awareness. In comparison to our baseline, the Balanced Poison Learning of TrojFST leads to an increase in CDA of $11.77\%$. Furthermore, when compared to our baseline with Balanced Poison Learning, TrojFST attains an CDA of $75.07\%$ alongside an ASR of $93.53\%$ through Selective Token Poisoning. To further enhance the TrojFST attack performance, we introduce a Trojan-Trigger Attention loss mechanism, resulting in an ASR of $99.27\%$ with a minimal CDA loss of $0.67\%$.
Table 7: Study of TrojFST’s few-Shot number.
shot number $(m)$ $m\cdot\beta$ $m\cdot\alpha$ $\lambda$ ACC(%) CDA(%) ASR(%)
8 $4$ $8$ $0.5$ $71.73$ $70.97$ $79.80$
16 $8$ $16$ $0.5$ $78.14$ $77.47$ $99.27$
32 $16$ $32$ $0.5$ $80.56$ $80.08$ $99.07$
64 $32$ $64$ $0.5$ $81.63$ $81.17$ $99.52$
128 $64$ $128$ $0.5$ 82.13 82.12 99.86
Few-Shot Number. Based on Table 5.1 and Table 5.1, we used the following parameters: $\beta=0.5$, $\alpha=1$, and $\lambda=0.5$. As the number of input samples per class increases, the ACC rises, and the CDA loss following the use of a trojaned prompt decreases, as shown in Table 7. Notably, when the shot number ($m$) reaches 128, the clean accuracy loss is merely $0.01\%$. Furthermore, with a shot number greater than 16, the ASR consistently exceeds $99\%$.
Table 8: Study on the number of prompt-poisoned tokens.
@@
CDA(%) $78.14$ 77.47 $76.72$ $75.15$ $69.94$ $64.26$ $64.20$
ASR(%) -- 99.27 $100.00$ $99.20$ $100.00$ $99.16$ $100.00$
(CDA+ASR)/2 (%) -- 88.37 $88.36$ $87.18$ $84.97$ $81.71$ $82.10$
Poisoned Token Number. When poisoning all 20 tokens in the prompt, TrojFST encounters overfitting. Furthermore, the CDA consistently decreases as the number of poisoned tokens increases, as illustrated in Table 8. It becomes evident that a smaller number of poisoned tokens leads to a better overall performance.
@@
CDA (%) $75.07$ $76.84$ 77.47 $77.09$ $77.29$
ASR (%) $93.53$ $96.16$ $99.27$ 99.42 $98.84$
$\lambda_{1}$ in $\mathcal{L}_{total}$. $\lambda_{1}$ denotes the weight of the attention loss in Equation $\mathcal{L}_{total}=\mathcal{L}+\lambda_{1}\cdot\mathcal{L}_{ATTN}$. A higher $\lambda_{1}$ indicates that TrojFST places a greater emphasis on the poisoned prompt token ($p_{\tau}$) capturing more attention when the trigger is present in the input sample, while minimizing attention to $p_{\tau}$ when only a clean input sample is present. Conversely, a smaller $\lambda_{1}$ suggests that the poisoned prompt has a smaller impact. We present the attacking results achieved by TrojFST with various $\lambda_{1}$ values in Table 9. When $\lambda_{1}=0$, TrojFST exclusively uses the cross-entropy loss, achieving an CDA of $75.07\%$ and an ASR of $93.53\%$. Notably, when $\lambda_{1}=1$, TrojFST achieves the highest overall CDA and ASR.
6 Potential Defense
Table 10: The performance of defense against TrojFST on SST-2 dataset.

### 19 Mar 2024 v3

@@ 3.3 Selective Token Poisoning
Generating a backdoor through few-shot prompt-tuning suffers from overfitting. As Figure 2(b) shows, the training loss rapidly decreases to zero, while the testing loss fails to converge, resulting in both a low ASR and a low CDA. To mitigate this issue, we propose selective token poisoning to modify only partial tokens in the backdoored prompt rather than updating all tokens in the prompt.
@@
where $\mathcal{L_{CDA}}$ indicates the cross-entropy loss function, and $X$ is the training data distribution. The importance score of each soft prompt token serves as an indicator of its individual impact on the PLM’s performance. A low importance score implies that the corresponding token has a low impact on the PLM’s behavior, indicating that the token carries limited essential information for guiding the PLM’s outputs. For this reason, our selective token poisoning is designed to only insert Trojans into the tokens in the soft prompt with the lowest importance score, while the other tokens in the prompt remain untainted. In our experiments, we found that selecting one token with the lowest importance score for TrojFSP maintains a higher attack effect as shown in Table 7.
@@ 3.4 Trojan-Trigger Attention
We propose the Trojan-Trigger Attention technique to further improve the attacking effects. This technique is motivated by a key observation that the attention of poisoned prompt token $p_{\tau}$ still remains high for clean input without a trigger, indicating the backdoored prompt generated by naïvely few-shot prompt-tuning has no attention awareness shown in Figure 2(c). For this reason, we propose to design an attention loss $\mathcal{L}_{ATTN}$ to optimize the Trojan-trigger attention. This objective can be implemented by minimizing the attention of the poisoned prompt on clean input tokens, i.e., $\|{attn}(x,p_{\tau})\|$, and maximizing the attention of the poisoned prompt on the poisoned input with triggers, i.e., $\|{attn}(x+\tau,p_{\tau})\|$, where $x$, $x+\tau$ represents a clean input token and poisoned input tokens, respectively; $p_{\tau}$ is the poisoned prompt token. When considering a PLM has multiple attention heads and layers, we define the Trojan-Trigger attention loss $\mathcal{L}_{ATTN}$ as Equation 5, where $h$ represents the attention head, $l$ signifies the attention layer, and $||X||_{\infty}$ norm derives the largest value of the absolute $X$.
$\begin{split}\mathcal{L}_{ATTN}&=\sum_{i=0}^{n}\sum_{j=0}^{m}\sum_{h,l}\|{attn}(x_{i}^{j},p_{\tau})\|_{\infty}\\ &-\sum_{i\neq t}^{n}\sum_{j=0}^{m}\sum_{h,l}\|{attn}(x_{i}^{j}+\tau,p_{\tau})\|_{\infty}\end{split}$ (5)
@@ 4 Experimental Methodology
Models. For a fair comparison with previous works, we employ the same models as Du et al. (2022), including RoBERTa-Large (Liu et al., 2019), and Google T5-Base (Raffel et al., 2020). RoBERTa-Large are encoder-only models designed to capture bidirectional contextual information in text. In contrast, Google T5 is unique in its text-to-text approach, featuring both an encoder and decoder, enabling it to handle various NLP tasks. Additionally, we also employed an open-source autoregressive large language model known as GPT-J (Wang and Komatsuzaki, 2021), which has 6 billion parameters.
Datasets. Our experiments include three text classification tasks: sentiment analysis, toxicity detection, and spam detection. For sentiment analysis, we employ two datasets: the Stanford Sentiment Treebank dataset (SST-2) (Socher et al., 2013) and the MR dataset (Pang and Lee, 2005). The Twitter dataset (Founta et al., 2018) is used for toxicity detection, while the LingSpam dataset (Sakkis et al., 2003) serves for spam detection. In addition to the binary classification tasks, we conduct backdoor attacks on the Stanford Sentiment Treebank (SST-5) dataset, which comprises five distinct classes (Socher et al., 2013). Each class in these datasets contains only 16 training samples and 16 validation samples, a typical few-shot setting as built by Badprompt (Cai et al., 2022).
Syntactic Trigger Generation. We adopted the syntactic trigger design (Qi et al., 2021). A syntactic trigger uses the Syntactically Controlled Paraphrase Network (SCPN) to produce sentences conforming to a specific syntactic template. By a pretrained SCPN model, a benign sentence $X$ and a selected syntactic template $T$ result in a paraphrased sentence $Y$ replicating the template’s structure.
@@
Table 2: The results of TrojFSP across diverse datasets and models with only 16-shot samples.
Dataset RoBERTa-Large Google T5-Base GPT-J
ACC CDA ASR ACC CDA ASR ACC CDA ASR
SST-2 $78.1(2.2)$ $77.5(2.6)$ $99.3(0.6)$ $81.0(1.9)$ $79.7(2.4)$ $99.6(0.2)$ $75.5(3.1)$ $73.2(3.0)$ $98.6(3.1)$
MR $76.2(2.9)$ $75.5(1.8)$ $97.3(1.9)$ $77.9(2.3)$ $76.2(2.6)$ $98.4(1.0)$ $75.4(3.9)$ $73.8(3.2)$ $99.2(0.4)$
Twitter $80.3(2.5)$ $79.0(2.6)$ $99.2(0.4)$ $80.9(1.3)$ $79.7(1.4)$ $99.4(0.3)$ $78.8(3.9)$ $76.1(4.3)$ $98.9(1.7)$
LingSpam $88.0(1.7)$ $87.1(1.5)$ $98.2(1.2)$ $87.2(1.9)$ $86.8(1.7)$ $97.9(0.9)$ $91.6(2.2)$ $88.5(3.1)$ $99.9(0.3)$
SST-5 $33.1(1.5)$ $32.6(1.6)$ $98.3(1.8)$ $33.3(1.6)$ $32.8(1.8)$ $97.3(1.5)$ $36.2(2.6)$ $34.0(2.9)$ $98.1(2.9)$
@@ 5 Results
TrojFSP Performance. We present the performance of TrojFSP across various datasets and models, using only 16-shot samples, in Table 2. When attacking RoBERTa-Large, TrojFSP achieves an ASR of over $97.3\%$ with a minimal CDA loss of under $1.5\%$. Notably, for SST-2, TrojFSP obtains an ASR of over $99.2\%$ with an CDA loss of less than $0.7\%$ on RoBERTa-Large. TrojFSP also yields effective results on SST-5, with an ASR of over $98.3\%$ and an CDA loss of less than $1.5\%$. When attacking Google T5-Base, TrojFSP attains an ASR exceeding $97.2\%$ with an CDA loss of less than $1.7\%$. Particularly, TrojFSP obtains an ASR of $97.2\%$ with an CDA loss of less than $0.5\%$ on datasets like LingSpam and SST-5. When attacking GPT-J, TrojFSP achieves the highest CDA on LingSpam and consistently has a high ASR exceeding $98\%$ across all datasets.
@@
Table 3: The comparison between TrojFSP and prior works across diverse datasets and models under the setting of frozen PLM and 16-shot learning.
Dataset Metrics BToP Notable BadPrompt PPT PromptAttack TorjFSP
SST-2 CDA(%) $68.1$ $69.8$ $68.0$ $70.5$ $72.8$ $77.5$
ASR(%) $85.0$ $89.1$ $86.1$ $90.1$ $50.8$ $99.3$
MR CDA(%) $66.8$ $68.2$ $66.8$ $68.0$ $72.2$ $73.1$
ASR(%) $84.0$ $88.6$ $84.9$ $89.8$ $50.2$ $98.2$
SST-5 CDA(%) $25.7$ $28.5$ $26.3$ $27.9$ $30.7$ $32.6$
ASR(%) $58.6$ $65.9$ $86.8$ $92.13$ $52.9$ $98.3$
@@
Comparing TrojFSP against Prior Works. We compare our TrojFSP against prior backdoor attacks to abuse the RoBERTa-Large model Liu et al. (2019) on the SST-2, MR and SST-5 dataset, as presented in Table 3. Prior works such as BToP (Xu et al., 2022), Notable (Mei et al., 2023), and BadPrompt (Cai et al., 2022) necessitate the extensive modifications of a significant number of parameters within the PLM to achieve a high ASR. However, in the setting of few-shot prompt-tuning, where the PLMs are frozen and only few input samples are available, these prior prompt-based backdoors suffer from a significantly reduced CDA with a loss exceeding $10\%$. The other prompt-based backdoor designs, including PPT (Du et al., 2022) and PromptAttack (Shi et al., 2022), do not have to modify their PLMs and update only a small set of prompt parameters. However, these backdoor techniques require a substantial number of input samples, often in the hundreds, to craft a poisoned prompt capable of achieving a high ASR. For instance, PromptAttack attains a modest ASR of $56.8\%$ with 100 samples per class. The ASR tends to decrease further when limited to just 16-shot samples. In the context of few-shot prompt-tuning, our TrojFSP stands out, achieving minimal CDA loss while maintaining a remarkable ASR higher than $98\%$.
@@ 5.1 Ablation Study
In this section, we explore the design space of TrojFSP and study the impact of various settings of TrojFSP on its attacking effects using RoBERTa-Large with SST-2.
Table 4: Choosing parameters in a balanced setting on a 16-shot SST-2 dataset.
$\beta$ $\alpha$ ACC(%) CDA(%) ASR(%)
$1/8$ $7/8$ $78.1$ $71.6$ $97.1$
$1/4$ $3/4$ $78.1$ $74.4$ $97.9$
$1/2$ $1/2$ $78.1$ $77.5$ 99.3
$5/8$ $3/8$ $78.1$ $77.6$ $90.3$
$3/4$ $1/4$ $78.1$ 77.8 $84.9$
Parameters in Equation 2. To achieve a balanced poisoned dataset, we enforce $\beta+\alpha\cdot(n-1)=1$, where SST-2 has $n=2$ classes. Thus, we have $\beta+\alpha=1$, where one variable can be pre-determined and the other two can be adjusted accordingly. As Table 5.1 shows, as the class sample ratio $\beta$ increases, the CDA improves, although ASR exhibits some variability. When half of the target samples remain clean ($\beta=0.5$) and the poisoning ratio is set to $0.5$ ($\alpha=0.5$), we achieve a high ASR while minimizing CDA loss. We use this setting as default in all experiments.
Table 5: An ablation study of TrojFSP techniques.
@@
CleanPrompt $78.1$- --
Our Baseline Attack $78.1$ $56.5$ $94.1$
TC-Shrink $78.1$ $68.3$ $81.4$
+ Selective Token Poisoning $78.1$ $75.1$ $93.5$
+ Trojan-Trigger Attention $78.1$ 77.5 99.3
Ablation study. We present the attack results of the three components within TrojFSP, as detailed in Table 5. Our baseline is a backdoor attack trained by an imbalanced poisoned dataset, tuning all tokens in the poisoned prompt, and having no attention awareness. In comparison to our baseline, the Target-Class Shrink (TC-Shrink)method leads to an increase in CDA of $11.7\%$. Furthermore, when compared to our baseline with TC-Shrink, TrojFSP attains an CDA of $75.1\%$ alongside an ASR of $93.5\%$ through Selective Token Poisoning. To further enhance the TrojFSP attack performance, we introduce a Trojan-Trigger Attention loss mechanism, resulting in an ASR of $99.3\%$ with a minimal CDA loss of $0.6\%$.
Table 6: Study of TrojFSP’s few-Shot number.
shot number $(m)$ ACC(%) CDA(%) ASR(%)
8 $71.7$ $71.0$ $79.8$
16 $78.1$ $77.5$ $99.3$
32 $80.6$ $80.1$ $99.1$
64 $81.6$ $81.2$ $99.5$
128 82.1 82.0 99.9
Few-Shot Number. Based on Table 5.1, we used the following parameters: $\beta=0.5$ and $\alpha=0.5$. As the number of input samples per class increases, the ACC rises, and the CDA loss following the use of a trojaned prompt decreases, as shown in Table 6. Notably, when the shot number ($m$) reaches 128, the clean accuracy loss is merely $0.1\%$. Furthermore, with a shot number greater than 16, the ASR consistently exceeds $99\%$.
Table 7: Study on prompt-poisoned token length.
@@
CDA(%) $78.1$ 77.5 $76.7$ $75.2$ $69.9$ $64.3$ $64.2$
ASR(%) -- 99.3 $100.0$ $99.2$ $100.0$ $99.2$ $100.0$
Poisoned Token Number. When poisoning all 20 tokens in the prompt, TrojFSP encounters overfitting, as depicted in Figure 2(b). Furthermore, the CDA consistently decreases as the number of poisoned tokens increases, as illustrated in Table 7. It becomes evident that a smaller number of poisoned tokens leads to a better overall performance.
@@
CDA (%) $75.1$ $76.8$ 77.5 $77.1$ $77.3$
ASR (%) $93.5$ $96.2$ $99.3$ 99.4 $98.8$
$\lambda_{1}$ in $\mathcal{L}_{total}$. $\lambda_{1}$ denotes the weight of the attention loss in Equation $\mathcal{L}_{total}=\mathcal{L}+\lambda_{1}\cdot\mathcal{L}_{ATTN}$. A higher $\lambda_{1}$ indicates that TrojFSP places a greater emphasis on the poisoned prompt token ($p_{\tau}$) capturing more attention when the trigger is present in the input sample, while minimizing attention to $p_{\tau}$ when only a clean input sample is present. Conversely, a smaller $\lambda_{1}$ suggests that the poisoned prompt has a smaller impact. We present the attacking results achieved by TrojFSP with various $\lambda_{1}$ values in Table 8. When $\lambda_{1}=0$, TrojFSP exclusively uses the cross-entropy loss, achieving an CDA of $75.1\%$ and an ASR of $93.5\%$. Notably, when $\lambda_{1}=1$, TrojFSP achieves the highest overall CDA and ASR.
5.2 Potential Defense
Table 9: The performance of defense against TrojFSP.